2. In a study on CHD (coronary heart disease), 200 individuals from each of the six age groups were taken, and presence or absence of CHD of each sample individual was recorded. The following data were obtained. CHD Age Present Absent [25,35) 10 190 35,45 23 177 [45,55) 71 (55,65 133 67 [65,75) 179 21 75,85 191 6 129

media%2F1f0%2F1f0de6c4-250d-45e5-a412-67
Answer #1

(a)

Overall incidence rate = Total CHD in each exposure group / Total observations

= (305 + 58) / (650 + 350) = 0.363

Incidence rate in Competitive exposure group = Total CHD in competitive exposure group / Total competitive observations

= 305 / 650 = 0.469

Incidence rate in NonCompetitive exposure group = Total CHD in noncompetitive exposure group / Total noncompetitive observations

= 58/350 = 0.166

b)

RR for Males

P0= probability of CHD for noncompetitive persons = 50/200 = 0.25

P1= probability of CHD for competitive persons = 300 / 600 = 0.5

Relative risk, risk ratio: RR = P1 / P0 = 0.5 / 0.25 = 2

RR for Females

P0= probability of CHD for noncompetitive persons = 8/150 = 0.053

P1= probability of CHD for competitive persons = 5 / 50 = 0.1

Relative risk, risk ratio: RR = P1 / P0 = 0.1 / 0.053 = 1.887

c)

Overall Risk ratio ,

P0= probability of CHD for noncompetitive persons = 58/350 = 0.166

P1= probability of CHD for competitive persons = 305 / 650 = 0.469

Relative risk, risk ratio: RR = P1 / P0 = 0.469 / 0.166 = 2.82

The risk ratio for the total, combined sample is RR = 2.82; this is sometimes referred to as the "crude" measure of association, because it is not adjusted for potential confounding factors. The risk ratios for the gender-stratified analysis are similar (RR = 2 and 1.887, respectively), but less than the crude risk ratio. This indicates that there was confounding by gender in the overall sample. As, the risk ratios for the gender-stratified analysis are similar (RR = 2 and 1.887, respectively), there is no effect modifier.

d)

As, RR is greater than 1, this means that competitive persons have a higher risk than non-competitive persons. Also, RR is near 2, means that the risk of CHD is doubled. In other words, the risk of CHD is increased by 100% or increased to 200%.